A car travels 100 miles in 2 hours and 30 minutes. Work out its average speed
Answer by rothauserc(4718)

rate * time = distance
rate = distance / time = 100 / 2.5 = (100*2) / 5 = 40 mph
checking our answer
40 * 5/2 = 100
100 = 100
our answer checks